Use linear approximation, i.e. the tangent line, to approximate 3.76 as follows:
Let f(x)=x6. The equation of the tangent line to f(x) at x=4 can be written in the form y=mx+b
where m is:

and where b is:

Using this, we find our approximation for 3.76 is
1
Expert's answer
2020-11-30T20:04:54-0500

Given "f(x)=x^6"

Tangent line to "f(m)" of "x=4"

at "x = 4 :

y = x^6;
y = 4^6 \implies" "(4, 4^6)"

Slope: "m(y) = y`= 6*5"

"m = 6*4^5"

"y - y_o = m(x-x_o)"

"y - 4^6 = 6*4^5[x-4]\\

\\y = x[6*4^5] - 24*4^5+46\\y=6*4^5[x] - 4^5*(24-4)"

"y = 6*4^5[x] - 20*4^5"

So compaing with y = m*x+c:

m = "6(4)^5;" c = "-20(4)^5"

To estimate value of "3.7^6" using linear approximation:

f(x) = f(a)+f`(a)(x-a)

so "f(3.7) = f(4) + f`(4)(3.7-4) = 4^6+6*4^5*(-0.3)"

"= 4^6 - 6*(0.3)*(4)^5 = 2252.8"

So, linear approximation to estimate "3.7^6" is 2252.8